OpenStudy (anonymous):

Let f(x) = x + 7 and g(x) = x – 2. What is f(x) • g(x)?

OpenStudy (john_es):

\[f(x)g(x)=(x+7)(x-2)=x^2+5x-14\]

OpenStudy (anonymous):

thanks, buuuuttttt I didn't mean to post this